OpenStudy (anonymous):

Part 1: Find the perimeter of a rectangular object which has a length of square root of 128 feet and a width of square root of 200 feet. (3 points) Part 2: Explain, in complete sentences, how you arrived at the answer and give the final solution in simplified radical form. (2 points) Part 3: What type of object in your home or school might this be? (1 point)

OpenStudy (yttrium):

Part 1. Perimeter of Rectangle. P = 2 (L+W) P = 2 (sqrt128+sqrt200) P = 2(8sqrt2 + 10sqrt2) P = 2(18sqrt2) P = 36sqrt2 feet.

OpenStudy (anonymous):

that i get i have figured out but what i can't figure out is what object that is. @Yttrium thank you for your help

OpenStudy (yttrium):

I bet you're the one to anser that. hahaha. :D